OpenStudy (anonymous):

Write the slope-intercept form of the equation that passes through (-1, -2) and is parallel to y=x-3

OpenStudy (anonymous):

@Twinz99

OpenStudy (anonymous):

@taylor12344

OpenStudy (anonymous):

y=x-1 Sorry it took a second to graph xD

OpenStudy (anonymous):

Slope intercept form is y=MX+B M is your slope and B is the Y intercept. You can have it intercepting at -1 with a slope of 1 and it's parallel and the equation is y=X-1